OpenStudy (anonymous):

If AD = 7 and AC = 28, calculate the length of line segment AB

OpenStudy (anonymous):

bad drawing sorry

OpenStudy (anonymous):



OpenStudy (anonymous):

C^2=A^2+B^2 Hence; 784=49+B^2 Doing the Math, we get the value of B as; 27 APPROXIMATELY; Which is basically the length of AB.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You welcome.